Session: Nov 2000 (session 4)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Huang Xinyong at Sun Nov 26 16:15:50 MET 2000) Success

Medicina (Mariotti at Sun Nov 26 18:15:35 MET 2000) Failure
We apologize, but due to electrical antenna problems we was able to observe only for the first four hours.

Effelsberg (A. Kraus at Mon Nov 27 09:08:38 MET 2000) Minor failures
Due to receiver problems and strong wind, we could not start observing before UT1235. Additonally, the scans between UT1500-1510 were lost, because a long slew was needed (cable wrap).

Metsahovi (Prisse Könönen at Tue Nov 28 08:59:51 MET 2000) Minor failures
We lost the first half of the session due to the need to restart the formatter, to reset times (the fs and the formatter) and, to redefine Tsys functions. The formatter and the time problems were due to the failure of the maser (in fri, 24th of Nov) which was caused by the broken maser room ventilation system. From 17:30 UTC the observation was ok. There were stable clouds.

Onsala (A. Polatidis at Tue Nov 28 10:08:56 MET 2000) Success
We observed only LCP, so expect signal on VC01 only.

Noto (C.Trigilio at Tue Nov 28 18:15:31 MET 2000) Success
Covered sky.

Jodrell_Bank (Alastair Gunn at Mon Dec 4 13:25:21 MET 2000) Success
No known problems.